Understand How to Use Specific Heat in Heat Loss/Gain Calculations Question Based on the following information, what is the specific heat? 400. g, q = 5.00 kJ, T, = 40.0, T-60.0 Report your answer with the correct number of significant figures. Provide your answer below: Jig degree C Content attribution Setup XFINITY.msi install flash player...mginstailflash player....dmg 23 20 F3 F4 F5 F6
Answer #1

We know, specific heat, Cs = q/m.∆T

We have q = 5.00 kJ

∆T = Tf - Ti = 60.0 - 40.0 = 20.0

m = 400 g

Now, putting all the values in the above equation,

Cs = 0.625 J/g degree C